Prove:
if 3 is not divisible by m * n, then 3 is not divisible by m and 3 is not divisible by n

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3816)


Proof by contradiction:

Assume that 3 is a factor of m. We'll show a contradiction arises from this assumption.
This means m = 3k for some integer k
This further leads to mn = 3k*n = 3(kn)
Showing that mn is a multiple of 3.
But this contradicts the fact that mn is not a multiple of 3.
Therefore, we must have m be a non-multiple of 3 as well.

Similar steps would apply to show that n must be a non-multiple of 3. This is one application of "without loss of generality" (WLOG) you can do.

Ultimately you should find that we get a contradiction if either m = 3k or n = 3p for integers k and p. Therefore, m and n cannot be multiples of 3.

Answer by ikleyn(52776)

.
Prove:
if 3 is not divisible by m * n, then 3 is not divisible by m and 3 is not divisible by n
~~~~~~~~~~~~~~~~~~~~


                    The posted statement is  INCORRECT.


One possible counter-example is  m = 5,  n = 1.

It is true that  3  is not divisible by  m*n = 5*1 = 5,  in this case - - - so the premise is  TRUE.

But the  "then"  statement is  FALSE:


    it is true that 3 is not divisible by 5, 

    but it is FALSE that 3 is not divisible by 1.


This counter-example  DISPROVE  your statement in the post.

---------------

Solved and  DISPROVED.         Answered and explained.


/////////////


From the first glance on your post it is clear that it is like a sock turned inside out,

and the correct statement must be different - but in this case,  I prefer do not discuss

my suspicions and instead disprove the wrong statement.